# Birational complexity and conic fibrations

### v1

@@ 2.5. Generalized canonical bundle formula
The following statement is known as the generalized canonical bundle formula. The proof is verbatim from [17, Theorem 1.5] by replacing [3, Theorem 1.4] with [30, Theorem 1.3].
@@
We will use an equivariant version of the previous statement. The following statement is known as the equivariant generalized canonical bundle formula for Fano type morphisms. The proof is verbatim from the proof of [35, Lemma 2.33] by replacing [17, Theorem 1.5] with Proposition 2.29.
@@ 2.7. Finite equivariant geometry
In this section, we prove some results regarding $G$-equivariant geometry. The first result is proved in [35, Theorem 5.1]
@@
The generalized pair $(X,(1-\epsilon)B,(1-\epsilon)\mathbf{M})$ is $G$-equivariant and klt. If $\epsilon>0$ is small enough, then the valuations in $\mathcal{V}$ are generalized non-canonical places of $(X,(1-\epsilon)B,(1-\epsilon)B)$. Hence, we may find a $G$-invariant effective divisor $\Gamma\sim_{\mathbb{Q}}(1-\epsilon)B+(1-\epsilon)\mathbf{M}_{X}$ such that $(X,\Gamma)$ is klt and every valuation in $\mathcal{V}$ is a non-canonical valuation of $(X,\Gamma)$ (see, e.g., [37, Lemma 3.7]). Thus, the statement follows from [34, Theorem 1]. ∎
@@ 2.8. Finite actions on conic fibrations
[…]
@@
(2.2) $1\rightarrow\mathbb{Z}_{2}\rightarrow G\rightarrow\mathbb{Z}_{m}\rightarrow 1$
@@ 2.9. Generalized pairs on toric varieties
Let $B$ be the reduced toric boundary of $X$ so $(X,B)$ is log Calabi–Yau toric pair. Let $\pi\colon Z\rightarrow X$ be a log resolution of $(X,B)$ for which the birational map $X\dashrightarrow Y$ extends to a morphism $\pi\colon Z\rightarrow Y$. Let $(Z,B_{Z})$ be the pair obtained by log pull-back of $(X,B)$ to $Z$. Note that $(Z,B_{Z})$ is a sub-log Calabi–Yau pair. Let $\Gamma_{Z}:=B_{Z}^{=1}$. The prime components of $\Gamma_{Z}$ are precisely the toric valuations extracted by $Z\rightarrow X$ (see Lemma 2.41). Fix $10<\epsilon<$ 1. Let $\Delta_{Z}$ be the boundary divisor obtained from $B_{Z}$ by increasing the coefficients of the prime components in ${\rm Ex}(\pi)\setminus\Gamma_{Z}$ to $\epsilon$. Then, we can write $\Delta_{Z}:=\Gamma_{Z}+\epsilon F_{Z}$ where $F_{Z}$ is reduced. The log pair $(Z,\Delta_{Z})$ is dlt and every prime component of $F_{Z}$ is exceptional over $X$. Therefore, the divisor $K_{Z}+\Gamma_{Z}+\epsilon F_{Z}$ is $\mathbb{Q}$-linearly equivalent to an effective divisor that is exceptional over $X$. The pair $(Z,\Gamma_{Z}+F_{Z})$ is dlt as $\Gamma_{Z}+F_{Z}$ has simple normal crossing support. So the divisor $F_{Z}$ contains no strata of $\Gamma_{Z}$. We run a $(K_{Z}+\Gamma_{Z}+\epsilon F_{Z})$-MMP over $X$ with scaling of an ample divisor. By Lemma 2.28, this MMP terminates in a model $W$ after contracting all the components of $F_{Z}$. Let $(W,\Gamma_{W})$ be the dlt pair induced on $W$. Note that $(W,\Gamma_{W})$ is a dlt modification of $(X,B)$. Hence, $W$ is a smooth projective toric variety. Let $q\colon Z\dashrightarrow W$ be the induced morphism. By monotonicity of log discrepancies, the subvariety $V:={\rm Ex}(q^{-1})$ does not contain any log canonical center of $(W,\Gamma_{W})$. Hence, $V$ does not contain any dlt center of $(W,\Gamma_{W})$. The dlt centers of $(W,\Gamma_{W})$ are precisely the toric strata of $W$. Thus, $V$ does not contain any toric strata of $W$. Note that $W\dashrightarrow Z$ is an isomorphism on the complement of $V$. Hence, the indeterminacy locus of $W\dashrightarrow Y$ is contained in $V$. So, the indeterminacy locus of $W\dashrightarrow Y$ does not contain any toric strata of $W$. Thus, it suffices to set $X^{\prime}:=W$. In our construction, $(W,\Gamma_{W})$ is a dlt toric pair so it is log smooth. In particular, $X^{\prime}$ is smooth. ∎
@@ 2.10. Algebraic torus actions
Let $X$ be a normal projective variety with a $\mathbb{G}_{m}$-action. Let $Y$ be the normalized Chow quotient of $X$ by the $\mathbb{G}_{m}$-action. By [2, Theorem 5.6], there is a divisorial fan $\mathcal{S}$ on $(Y,\mathbb{Q})$ for which there is a $\mathbb{G}_{m}$-equivariant isomorphism $X(\mathcal{S})\simeq X$. The tail-fan of $\mathcal{S}$ is $\Sigma=\{\rho_{-},\{0\},\rho_{+}\}$, where $\rho_{-}:=\mathbb{Q}_{\leq 0}$ and $\rho_{+}:=\mathbb{Q}_{\geq 0}$. Ther rays $\rho_{-}$ and $\rho_{+}$ induce two $\mathbb{G}_{m}$-invariant divisorial valuations $E_{-}$ and $E_{+}$ over $X$. We call $E_{-}$ the divisorial sink and $E_{+}$ divisorial source of the $\mathbb{G}_{m}$-action, respectively.
Let $\widetilde{X}\rightarrow X$ be a $\mathbb{G}_{m}$-equivariant projective birational morphism for which there is a $\mathbb{G}_{m}$-quotient $\widetilde{X}\rightarrow Y$. The center of the divisorial sink and the divisorial source of $\mathbb{G}_{m}$ on $\widetilde{X}$ are the two $\mathbb{G}_{m}$-invariant prime divisors that dominate $Y$. These are precisely the only two prime divisors on $\widetilde{X}$ on which $\mathbb{G}_{m}$ acts as the identity (see, e.g., [1, Theorem 10.1]). Hence, the sink and the source of the $\mathbb{G}_{m}$-action on $X$ are the centers of the divisorial sink and the divisorial source, respectively.
@@
By [42, Theorem 1], we have a $\mathbb{G}_{m}$-equivariant isomorphism
@@ 3. Finite actions on Fano fibrations
By assumption, a general fiber $X_{z}$ is and $\epsilon$-lc Fano variety and $N(K_{X_{z}}+B_{z}+\mathbf{M}_{z,X_{z}})\sim 0$. In particular, the pairs $(X_{z},B_{z})$ as in the statement are log bounded in terms of $n,N$, and $\epsilon$ (see, e.g., [5]). By [34, Theorem 4.1], there is a constant $m_{0}:=m_{0}(n,N,\epsilon)$ satisfying the following. Whenever $\mathbb{Z}_{\ell}\leqslant{\rm Aut}(X,B)$ acts fiberwise over $Z$ and $\ell\geq m_{0}$, we can find a subgroup $C\leqslant\mathbb{Z}_{\ell}$ of index at most $\frac{m_{0}}{2}$ such that $C\leqslant\mathbb{T}_{z}\leqslant{\rm Aut}^{0}(X_{z},B_{z})$ for a general closed point $z\in Z$. Here, $\mathbb{T}_{z}$ is a maximal torus of ${\rm Aut}^{0}(X_{z},B_{z})$. Let $\eta$ be the generic point of $Z$ and $\bar{\eta}$ be the geometric generic point. We write $(X_{\eta},B_{\eta})$ and $(X_{\bar{\eta}},B_{\bar{\eta}})$ for the log pair induced on the generic fiber and the geometric generic fiber, respectively. By [41, Lemma 2.1], we conclude that $C\leqslant\mathbb{T}_{\bar{\eta}}\leqslant{\rm Aut}(X_{\bar{\eta}},B_{\bar{\eta}})$, where $\mathbb{T}_{\bar{\eta}}$ is a split algebraic torus over $\bar{\eta}$ of rank $r\leq n$. We write $\mathbb{T}_{\eta}$ for the corresponding algebraic torus over $\eta$. In particular, up to shrinking $Z$ and assuming that the base is affine, we have $C\leqslant\mathbb{T}_{Z}\leqslant{\rm Aut}_{Z}(X,B)$ where $\mathbb{T}_{Z}$ is an algebraic torus over $Z$.
@@
${\bar{X}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{q}$${\bar{X}/C\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${U_{W}}$
@@
$1\rightarrow\mathbb{T}_{Z^{\prime},0}\rightarrow H_{0}\rightarrow G\rightarrow 1$
@@
Let $\mathbb{T}_{\bar{\eta},0}$ be the one-dimensional torus induced by $\mathbb{T}_{Z^{\prime},0}$ on the geometric generic fiber. So, we have $\mathbb{T}_{\bar{\eta},0}\leqslant{\rm Aut}(X_{\bar{\eta}},B_{\bar{\eta}})$. For each $m\geq 1$, we let $H_{0,m}$ be the subgroup of $H_{0}$ generated by $\mathbb{Z}_{m}$ and $G$. Note that $(X^{\prime},B^{\prime})$ is $H_{0,m}$-invariant and $\mathbb{Q}$-complemented over $Z$. By [35, Lemma 2.14], for each $m\geq 1$, we can shrink $Z$ and construct an effective divisor $\Gamma_{m}$ for which:
@@
${Y\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${Y^{\prime}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{/G}$${X}$${X^{\prime}\ignorespaces\ignorespaces\ignorespaces\ignorespaces.}$$\scriptstyle{/G}$
@@ 4. Finite abelian actions and conic fibrations
[…]
@@
By [6, Theorem 1.6], there exists a finite subset $\mathcal{F}\subseteq\Lambda$ such that the coefficients of $B$ belong to $\mathcal{F}$ and the coefficients of $\mathbf{M}$ belong to $\mathcal{F}$ in a quotient where it descends. Hence, by [17, Theorem 1.2], we conclude that $I(K_{X}+B+\mathbf{M}_{X})\sim 0$ for some $I$ that only depends on $n$ and $\mathcal{F}$. In particular, the positive integers $I$ only depends on $n$ and $\Lambda$.
@@
Assume that ${\rm reg}(X_{z},B_{z},\mathbf{M}_{z})\geq 1$. This condition implies that the dimension of $X_{z}$ is at least $2$. Let $U\subseteq Z$ be an open set for which there is an $A$-invariant $N$-complement $\Gamma_{U}$ of $(X_{U},B_{U})$ over $Z$. Here, we may choose $N$ so it only depends on $n$ and $\Lambda$ (see [17, Theorem 1.2]). Further, the pair $(X_{U},B_{U})$ is the restriction of $(X,B)$ to the preimage of $U$. We may assume that ${\rm reg}(X_{z},\Gamma_{z})\geq{\rm reg}(X_{z},B_{z},\mathbf{M}_{z})\geq 1$. Let $m_{2}:=m_{2}(n,N,I^{-1})$ be as in Proposition 2.35. Assume that $|A_{f}|>m_{1}(n,I,\epsilon):=m_{2}^{n}$. Since $A_{f}$ is a finite abelian group of rank at most $n$, we conclude that there is a cyclic subgroup $C\leqslant A_{f}$ of order at least $m_{2}$. Note that $C\leqslant{\rm Aut}(X_{U},\Gamma_{U})$ is acting fiberwise over $U$, i.e., we have $C\leqslant{\rm Aut}_{U}(X_{U},\Gamma_{U})$. By Theorem 3.3, there exists a subgroup $C_{0}\leqslant C$ of index at most $\frac{m_{2}}{2}$ satisfying the following. There exist:
@@
[…]
By induction, we know that there exists a subset $I\subseteq\{1,\dots,r+1\}$ such that $|I|=k-1$ and each $\psi_{i}$ with $i\in I$ is a conic fibration. Furthermore, if $i\in I$, $A_{Z,i}$ is the homomorphic image of $A_{Z,0}$ acting on $Z_{i}$, and $(Z_{i},B_{Z_{i}},\mathbf{N}_{i})$ is induced by the generalized canonical bundle formula on $Z_{i}$, then $\lfloor B_{Z_{i}}\rfloor$ admits two disjoint $A_{Z,i}$-invariant prime components that are horizontal over $Z_{i+1}$. Let $A_{0}\leqslant A$ be the preimage of $A_{Z,0}$ in $A$. Hence $A_{0}$ has index at most $m_{0}(\dim Z,\Omega)/2$ in $A$. By Proposition 2.31, there exists an $A_{0}$-equivariant crepant birational model $(X^{\prime},B^{\prime},\mathbf{M})$ that admits an equivariant conic fibration $\phi_{0}$ onto $Z^{\prime}$. By Proposition 2.31.(5), we have the following commutative diagram:
[…]
@@
[…]
@@
[…]

### v2

@@ 2.5. Generalized canonical bundle formula
The following statement is known as the generalized canonical bundle formula. The proof is verbatim from [17, Theorem 1.5] by replacing [4, Theorem 1.4] with [30, Theorem 1.3].
@@
We will use an equivariant version of the previous statement. The following statement is known as the equivariant generalized canonical bundle formula for Fano type morphisms. The proof is verbatim from the proof of [38, Lemma 2.33] by replacing [17, Theorem 1.5] with Proposition 2.29.
@@ 2.7. Finite equivariant geometry
In this section, we prove some results regarding $G$-equivariant geometry. The first result is proved in [38, Theorem 5.1]
@@
The generalized pair $(X,(1-\epsilon)B,(1-\epsilon)\mathbf{M})$ is $G$-equivariant and klt. If $\epsilon>0$ is small enough, then the valuations in $\mathcal{V}$ are generalized non-canonical places of $(X,(1-\epsilon)B,(1-\epsilon)B)$. Hence, we may find a $G$-invariant effective divisor $\Gamma\sim_{\mathbb{Q}}(1-\epsilon)B+(1-\epsilon)\mathbf{M}_{X}$ such that $(X,\Gamma)$ is klt and every valuation in $\mathcal{V}$ is a non-canonical valuation of $(X,\Gamma)$ (see, e.g., [32, Lemma 3.7]). Thus, the statement follows from [37, Theorem 1]. ∎
@@ 2.8. Finite actions on conic fibrations
${(X,B)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(X^{\prime},B^{\prime})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{/\mathbb{Z}_{2}}$${Z}$${Z^{\prime}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{/\mathbb{Z}_{2}}$
@@
(2.2) $1\rightarrow\mathbb{Z}_{2}\rightarrow G\rightarrow\mathbb{Z}_{m}\rightarrow 1.$
@@ 2.9. Generalized pairs on toric varieties
Let $B$ be the reduced toric boundary of $X$ so $(X,B)$ is log Calabi–Yau toric pair. Let $\pi\colon Z\rightarrow X$ be a log resolution of $(X,B)$ for which the birational map $X\dashrightarrow Y$ extends to a morphism $\pi\colon Z\rightarrow Y$. Let $(Z,B_{Z})$ be the pair obtained by log pull-back of $(X,B)$ to $Z$. Note that $(Z,B_{Z})$ is a sub-log Calabi–Yau pair. Let $\Gamma_{Z}:=B_{Z}^{=1}$. The prime components of $\Gamma_{Z}$ are precisely the toric valuations extracted by $Z\rightarrow X$ (see Lemma 2.41). Fix $0<\epsilon<1$. Let $\Delta_{Z}$ be the boundary divisor obtained from $B_{Z}$ by increasing the coefficients of the prime components in ${\rm Ex}(\pi)\setminus\Gamma_{Z}$ to $\epsilon$. Then, we can write $\Delta_{Z}:=\Gamma_{Z}+\epsilon F_{Z}$ where $F_{Z}$ is reduced. The log pair $(Z,\Delta_{Z})$ is dlt and every prime component of $F_{Z}$ is exceptional over $X$. Therefore, the divisor $K_{Z}+\Gamma_{Z}+\epsilon F_{Z}$ is $\mathbb{Q}$-linearly equivalent to an effective divisor that is exceptional over $X$. The pair $(Z,\Gamma_{Z}+F_{Z})$ is dlt as $\Gamma_{Z}+F_{Z}$ has simple normal crossing support. So the divisor $F_{Z}$ contains no strata of $\Gamma_{Z}$. We run a $(K_{Z}+\Gamma_{Z}+\epsilon F_{Z})$-MMP over $X$ with scaling of an ample divisor. By Lemma 2.28, this MMP terminates in a model $W$ after contracting all the components of $F_{Z}$. Let $(W,\Gamma_{W})$ be the dlt pair induced on $W$. Note that $(W,\Gamma_{W})$ is a dlt modification of $(X,B)$. Hence, $W$ is a smooth projective toric variety. Let $q\colon Z\dashrightarrow W$ be the induced morphism. By monotonicity of log discrepancies, the subvariety $V:={\rm Ex}(q^{-1})$ does not contain any log canonical center of $(W,\Gamma_{W})$. Hence, $V$ does not contain any dlt center of $(W,\Gamma_{W})$. The dlt centers of $(W,\Gamma_{W})$ are precisely the toric strata of $W$. Thus, $V$ does not contain any toric strata of $W$. Note that $W\dashrightarrow Z$ is an isomorphism on the complement of $V$. Hence, the indeterminacy locus of $W\dashrightarrow Y$ is contained in $V$. So, the indeterminacy locus of $W\dashrightarrow Y$ does not contain any toric strata of $W$. Thus, it suffices to set $X^{\prime}:=W$. In our construction, $(W,\Gamma_{W})$ is a dlt toric pair so it is log smooth. In particular, $X^{\prime}$ is smooth. ∎
@@ 2.10. Algebraic torus actions
Let $X$ be a normal projective variety with a $\mathbb{G}_{m}$-action. Let $Y$ be the normalized Chow quotient of $X$ by the $\mathbb{G}_{m}$-action. By [1, Theorem 5.6], there is a divisorial fan $\mathcal{S}$ on $(Y,\mathbb{Q})$ for which there is a $\mathbb{G}_{m}$-equivariant isomorphism $X(\mathcal{S})\simeq X$. The tail-fan of $\mathcal{S}$ is $\Sigma=\{\rho_{-},\{0\},\rho_{+}\}$, where $\rho_{-}:=\mathbb{Q}_{\leq 0}$ and $\rho_{+}:=\mathbb{Q}_{\geq 0}$. Ther rays $\rho_{-}$ and $\rho_{+}$ induce two $\mathbb{G}_{m}$-invariant divisorial valuations $E_{-}$ and $E_{+}$ over $X$. We call $E_{-}$ the divisorial sink and $E_{+}$ divisorial source of the $\mathbb{G}_{m}$-action, respectively.
Let $\widetilde{X}\rightarrow X$ be a $\mathbb{G}_{m}$-equivariant projective birational morphism for which there is a $\mathbb{G}_{m}$-quotient $\widetilde{X}\rightarrow Y$. The center of the divisorial sink and the divisorial source of $\mathbb{G}_{m}$ on $\widetilde{X}$ are the two $\mathbb{G}_{m}$-invariant prime divisors that dominate $Y$. These are precisely the only two prime divisors on $\widetilde{X}$ on which $\mathbb{G}_{m}$ acts as the identity (see, e.g., [2, Theorem 10.1]). Hence, the sink and the source of the $\mathbb{G}_{m}$-action on $X$ are the centers of the divisorial sink and the divisorial source, respectively.
@@
By [44, Theorem 1], we have a $\mathbb{G}_{m}$-equivariant isomorphism
@@ 3. Finite actions on Fano fibrations
By assumption, a general fiber $X_{z}$ is and $\epsilon$-lc Fano variety and $N(K_{X_{z}}+B_{z}+\mathbf{M}_{z,X_{z}})\sim 0$. In particular, the pairs $(X_{z},B_{z})$ as in the statement are log bounded in terms of $n,N$, and $\epsilon$ (see, e.g., [6]). By [37, Theorem 4.1], there is a constant $m_{0}:=m_{0}(n,N,\epsilon)$ satisfying the following. Whenever $\mathbb{Z}_{\ell}\leqslant{\rm Aut}(X,B)$ acts fiberwise over $Z$ and $\ell\geq m_{0}$, we can find a subgroup $C\leqslant\mathbb{Z}_{\ell}$ of index at most $\frac{m_{0}}{2}$ such that $C\leqslant\mathbb{T}_{z}\leqslant{\rm Aut}^{0}(X_{z},B_{z})$ for a general closed point $z\in Z$. Here, $\mathbb{T}_{z}$ is a maximal torus of ${\rm Aut}^{0}(X_{z},B_{z})$. Let $\eta$ be the generic point of $Z$ and $\bar{\eta}$ be the geometric generic point. We write $(X_{\eta},B_{\eta})$ and $(X_{\bar{\eta}},B_{\bar{\eta}})$ for the log pair induced on the generic fiber and the geometric generic fiber, respectively. By [43, Lemma 2.1], we conclude that $C\leqslant\mathbb{T}_{\bar{\eta}}\leqslant{\rm Aut}(X_{\bar{\eta}},B_{\bar{\eta}})$, where $\mathbb{T}_{\bar{\eta}}$ is a split algebraic torus over $\bar{\eta}$ of rank $r\leq n$. We write $\mathbb{T}_{\eta}$ for the corresponding algebraic torus over $\eta$. In particular, up to shrinking $Z$ and assuming that the base is affine, we have $C\leqslant\mathbb{T}_{Z}\leqslant{\rm Aut}_{Z}(X,B)$ where $\mathbb{T}_{Z}$ is an algebraic torus over $Z$.
@@
[…]
@@
$1\rightarrow\mathbb{T}_{Z^{\prime},0}\rightarrow H_{0}\rightarrow G\rightarrow 1.$
@@
Let $\mathbb{T}_{\bar{\eta},0}$ be the one-dimensional torus induced by $\mathbb{T}_{Z^{\prime},0}$ on the geometric generic fiber. So, we have $\mathbb{T}_{\bar{\eta},0}\leqslant{\rm Aut}(X_{\bar{\eta}},B_{\bar{\eta}})$. For each $m\geq 1$, we let $H_{0,m}$ be the subgroup of $H_{0}$ generated by $\mathbb{Z}_{m}$ and $G$. Note that $(X^{\prime},B^{\prime})$ is $H_{0,m}$-invariant and $\mathbb{Q}$-complemented over $Z$. By [38, Lemma 2.14], for each $m\geq 1$, we can shrink $Z$ and construct an effective divisor $\Gamma_{m}$ for which:
@@
[…]
@@ 4. Finite abelian actions and conic fibrations
    ${X^{\prime}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\psi_{0}}$${X_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\psi_{1}}\textstyle{\dots\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\psi_{r-1}}$${X_{r}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\psi_{r}}$${W}$
@@
By [3, Theorem 1.6], there exists a finite subset $\mathcal{F}\subseteq\Lambda$ such that the coefficients of $B$ belong to $\mathcal{F}$ and the coefficients of $\mathbf{M}$ belong to $\mathcal{F}$ in a quotient where it descends. Hence, by [17, Theorem 1.2], we conclude that $I(K_{X}+B+\mathbf{M}_{X})\sim 0$ for some $I$ that only depends on $n$ and $\mathcal{F}$. In particular, the positive integers $I$ only depends on $n$ and $\Lambda$.
@@
Assume that ${\rm reg}(X_{z},B_{z},\mathbf{M}_{z})\geq 1$. This condition implies that the dimension of $X_{z}$ is at least $2$. Let $U\subseteq Z$ be an open set for which there is an $A$-invariant NN-complement $\Gamma_{U}$ of $(X_{U},B_{U})$ over $Z$. Here, we may choose $N$ so it only depends on $n$ and $\Lambda$ (see [17, Theorem 1.2]). Further, the pair $(X_{U},B_{U})$ is the restriction of $(X,B)$ to the preimage of $U$. We may assume that ${\rm reg}(X_{z},\Gamma_{z})\geq{\rm reg}(X_{z},B_{z},\mathbf{M}_{z})\geq 1$. Let $m_{2}:=m_{2}(n,N,I^{-1})$ be as in Proposition 2.35. Assume that $|A_{f}|>m_{1}(n,I,\epsilon):=m_{2}^{n}$. Since $A_{f}$ is a finite abelian group of rank at most $n$, we conclude that there is a cyclic subgroup $C\leqslant A_{f}$ of order at least $m_{2}$. Note that $C\leqslant{\rm Aut}(X_{U},\Gamma_{U})$ is acting fiberwise over $U$, i.e., we have $C\leqslant{\rm Aut}_{U}(X_{U},\Gamma_{U})$. By Theorem 3.3, there exists a subgroup $C_{0}\leqslant C$ of index at most $\frac{m_{2}}{2}$ satisfying the following. There exist:
@@
${Z^{\prime}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{1}}$${Z_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{2}}\textstyle{\dots\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{r}}$${Z_{r}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{r+1}}$${W.}$
By induction, we know that there exists a subset $I\subseteq\{1,\dots,r+1\}$ such that |I|=k1|I|=k-1 and each $\psi_{i}$ with $i\in I$ is a conic fibration. Furthermore, if $i\in I$, $A_{Z,i}$ is the homomorphic image of $A_{Z,0}$ acting on $Z_{i}$, and $(Z_{i},B_{Z_{i}},\mathbf{N}_{i})$ is induced by the generalized canonical bundle formula on $Z_{i}$, then $\lfloor B_{Z_{i}}\rfloor$ admits two disjoint $A_{Z,i}$-invariant prime components that are horizontal over $Z_{i+1}$. Let $A_{0}\leqslant A$ be the preimage of $A_{Z,0}$ in $A$. Hence $A_{0}$ has index at most $m_{0}(\dim Z,\Omega)/2$ in $A$. By Proposition 2.31, there exists an $A_{0}$-equivariant crepant birational model $(X^{\prime},B^{\prime},\mathbf{M})$ that admits an equivariant conic fibration $\phi_{0}$ onto $Z^{\prime}$. By Proposition 2.31.(5), we have the following commutative diagram:
${(X,B,\mathbf{M})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(X^{\prime},B^{\prime},\mathbf{M})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{0}}$${(Z,B_{Z},\mathbf{N})}$${(Z^{\prime},B_{Z^{\prime}},\mathbf{N})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{1}}\textstyle{\vdots\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{r+1}}$${W}$
@@
${Z^{\prime}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{1}}$${Z_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{2}}\textstyle{\dots\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{r}}$${Z_{r}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{r+1}}$${W.}$
@@
${(X,B,\mathbf{M})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(X^{\prime},B^{\prime},\mathbf{M})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{0}}$${(Z,B_{Z},\mathbf{N})}$${(Z^{\prime},B_{Z^{\prime}},\mathbf{N})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{1}}\textstyle{\vdots\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{r+1}}$${W}$